OpenStudy (perl):

im a little confused about the zero product property suppose that you are asked to solve x^2 + 5x + 6 = 0 so we factor (x+3)(x+2) = 0 , and the zero product theorem says that x = -3 or x = -2. but i dont see why this is a solution to the original polynomial. and why isnt it x = -3 and x = -2, why or?

OpenStudy (perl):

so i am asking two different questions

jimthompson5910 (jim_thompson5910):

because 'and' means that x is BOTH values AT THE SAME TIME, which is impossible the 'or' means that x is either one value or the other (but not at the same time)

jimthompson5910 (jim_thompson5910):

not both at the same time*

OpenStudy (perl):

ok, and why are these the solutions to the original polynomial ?

jimthompson5910 (jim_thompson5910):

if you replace x with either -3 OR -4, and evaluate/simplify, you'll get 0

OpenStudy (anonymous):

But the "or" doesn't mean that they aren't both solutions, right? It's just that x can't be both values at the same time, right?

OpenStudy (perl):

why does factoring the polynomial and saying x = a or x = b, how come that solves the original equation

jimthompson5910 (jim_thompson5910):

the "or" means that either one is a solution, but you use it over "and" to avoid saying x is both values at the same time

jimthompson5910 (jim_thompson5910):

so the "or" is implying that both are the solutions (just not at the same time)

OpenStudy (perl):

why is x = a or x = b solutions to x^2 - (a+b)x + ab = 0 ?

OpenStudy (amistre64):

the poly factors into the product of to binoms

OpenStudy (amistre64):

*two

OpenStudy (perl):

but the funny thing is, if x = 2 or x = 3, its possible that x is not equal to 2 , but x = 3.

OpenStudy (amistre64):

funny thing?

jimthompson5910 (jim_thompson5910):

x^2 - (a+b)x + ab = 0 a^2 - (a+b)a + ab = 0 ... Replace all 'x' terms with 'a' (since x = a is one solution) a^2 - (a^2 + ab) + ab = 0 a^2 - a^2 - ab + ab = 0 0 = 0 So this shows that x = a is a solution to x^2 - (a+b)x + ab = 0

jimthompson5910 (jim_thompson5910):

well if x = 3, then x can't equal 2 (x can only equal one thing)

OpenStudy (amistre64):

x is any value from -inf to inf; it just so happens that when x=-2, the function equals zero; and then when x=3, the function equals zero

OpenStudy (perl):

yes, but why in general does factoring work for the original polynomial

jimthompson5910 (jim_thompson5910):

but x can certainly equal 2

OpenStudy (perl):

yes but it is 'or' so x = -2 OR x = -3, so its possible that x is not equal to -3 ... ok i think im missing some theorem here . is this a factor theorem,

OpenStudy (amistre64):

weve been multiplying things since what, the 3rd grade? x+3 x+2 ----- x^2+3x +2x+6 ----------- x^2+5x+6

OpenStudy (perl):

if x is not equal to -2, then x is equal to -3. if x is not equal to -2, then x is equal to -3. x cannot be both equal to -2 and -3. ok but im missing something, how do we know this solves the original polynomial without actually substituting (long way)

OpenStudy (perl):

the original equation is x^2 + 5x + 6 = 0, so im asking, why does solving the factored form of the polynomial solve the original polynomial equation

OpenStudy (amistre64):

the zero factor thrm ... also heard it called the zero product rule; is what happens when we multiply a term by zero, we get a zero back

OpenStudy (perl):

ok, you get a zero back. but how does that show the original polynomial (expanded) is a solution

OpenStudy (amistre64):

when does ab = 0? when either a or b (or both) are zero 0*b = 0 a*0 = 0 0*0 = 0 we learned this form zero multiplication tables

OpenStudy (perl):

but these are different expressions. one is the polynomial x^2 + 5x + 6, the other is (x+2)(x+3), they are very different

OpenStudy (amistre64):

when does x^2 + 5x + 6 = 0 ?? is the same as saying when does: (x+2)(x+3) = 0 ; well, when x+2 = 0 or when x+3 = 0

OpenStudy (perl):

why does the zeroes of (x+2)(x+3) make x^2+5x+6 also zero?

OpenStudy (amistre64):

why do you insist that they are different?

OpenStudy (amistre64):

becasue when you multiply out the products you get the poly .... as srawned out in a previous post

OpenStudy (perl):

but , i'm not comfortable with your logic you prove that x^2+5x+6 = 0 when x +2 = 0 or x + 3 = 0. so that means x could equal to 50 or -3 (since it is an 'or' statement )

OpenStudy (perl):

the truth value of the statement ' x = -3 or x = -2 ' is equivalent to the truth value of ' x = 50 or x = -2' . since they are both true when x = -2

OpenStudy (amistre64):

would you rather do this with math instead of making up nonmath things to try to confound the issue?

OpenStudy (perl):

an 'or' statement means that only one of the parts has to be true. im not making up non -math, i dont like your tone

OpenStudy (amistre64):

what does (x+2)(x+3) equal?

OpenStudy (perl):

this is basic logic, you should study it

OpenStudy (amistre64):

your not applying logic either

OpenStudy (perl):

yes i know it is equal to x^2 + 5x + 6. im asking about the 'or' statement

OpenStudy (perl):

yes i am, 'p or q' is true when p is true.

OpenStudy (perl):

'p or q' is true when p is true and q is false

OpenStudy (amistre64):

the or statement says that when x= one or the other, then we get a zero; you are trying to say that x can be more than one value at a time

OpenStudy (perl):

so ' x = -3 or x = 50 ' is true when x = -3 , so is ' x = -3 or x = -2 '

OpenStudy (perl):

no im saying one is enough , one solution

OpenStudy (perl):

are these steps reversible, maybe thats what im looking for

OpenStudy (perl):

we want two solutions, but the 'or' only guarantees one solution is true

OpenStudy (amistre64):

you might be confusing the english use of or as an exclusive or instead of that other one

OpenStudy (amistre64):

x is either one or the other in order for this to be zero

OpenStudy (perl):

' p or q ' is true when p is true, q is false p is false, q is true p is true, q is true

OpenStudy (perl):

we cannot infer from ' john is 8 years old or john live in california' that both are true. we only know at least ONE is true

OpenStudy (perl):

so saying ' x = -2 or x = -3' only tells me ONE of these are the solutions.

OpenStudy (perl):

but i dont know which one ( we can never know )

OpenStudy (perl):

strictly from logic :)

OpenStudy (amistre64):

and you also seem to be stretching the logic; 8 years and california are not both in the same domain John is 8 years old or John is 32 years old

OpenStudy (perl):

so again we cannot infer from 'p or q' , you cannot infer p is true or that q is true.

OpenStudy (perl):

i wanted to use different propositions

OpenStudy (perl):

'john is 8 years old or mike is 24 ' , you cannot infer that john is 8 , and you cannot infer that mike is 24

OpenStudy (amistre64):

but the original setup doesnt use different propositions; you have to construct the analogy to match the original or else its useless

OpenStudy (amistre64):

we dont have 2 different variables; no mike and john; just john

OpenStudy (perl):

well so far we have an 'or' statement. so we cannot infer which one is the solution, do you know what i mean. all we can say is that one of them is the solution

OpenStudy (perl):

ok, 'john is 8 years old or john is 24 '

OpenStudy (amistre64):

your misuse of logic aside, the original statement is sound :)

OpenStudy (perl):

we dont know which of those is true, all we know is that if the 'or' statement is true, one of them is true

OpenStudy (amistre64):

Let say John is a function defined as when John is 8 or when John is 32; he lives at home, otherwise he lives someplace else that isnt home when does john live at home?

OpenStudy (perl):

hmm

OpenStudy (perl):

what does living at home have to do with age

OpenStudy (amistre64):

what does that matter, its a rule that defines a function

OpenStudy (perl):

im not sure

OpenStudy (amistre64):

as we journey along the poly curve, there are only specific values of x that we come across that make the function zero

OpenStudy (perl):

but all im saying is, when you have an 'or' statement, you cannot know which of them is true . p or q is true , we dont know if p is true or q is true

OpenStudy (amistre64):

there are only certain times in johns live that he lives at home

OpenStudy (perl):

, all we can say is at least one of p or q is true (possibly both, depending on the problem)

jimthompson5910 (jim_thompson5910):

If you want to do this from a strict logical standpoint, then you might do something like this There are some numbers that satisfy x^2 + 5x + 6 = 0 There are some numbers that satisfy (x+3)(x+2) = 0 There are some numbers that satisfy x+3 = 0 or x+2 = 0 There are some numbers that satisfy x = -3 or x = -2 Keep in mind that (p v q) <==> (q v p), so we can easily say "There are some numbers that satisfy x = -2 or x = -3" BUT You cannot say (p v q) <==> (p v r), which is why you can't go from "There are some numbers that satisfy x = -3 or x = -2" to "There are some numbers that satisfy x = -3 or x = -50"

OpenStudy (perl):

but im saying, suppose we label p: ' x = -2' q: ' x = -3' , and that we say 'p or q' is true.

OpenStudy (perl):

we don't know which one is true , whether x is equal to -2 or whether x = -3 ,

OpenStudy (amistre64):

if (x=a or x=b), then (f(x)=0) is the statement to focus on; this gives the proper form for p and q

jimthompson5910 (jim_thompson5910):

both are true because both are part of the solution set

OpenStudy (perl):

hmm, but an 'or' statement only guarantees one of them is true. and we dont know which

jimthompson5910 (jim_thompson5910):

if you solve x^2 + 5x + 6 = 0, you get the solution set {-2, -3} or you can start with the idea that the solution set is {-2, -3} to get k(x^2 + 5x + 6) = 0 for some constant k

OpenStudy (amistre64):

there are 2 or statments that are used in english

OpenStudy (perl):

if i say it is true that ' john lives in germany or john is 28' we dont know which one is true. john could live in australia and be 28 , and it still be true

OpenStudy (perl):

or john could be 12 and live in germany, and it is still true

OpenStudy (amistre64):

but i might be stretching the xor assessment :)

OpenStudy (amistre64):

when john is 12, does he live at home? which is what the function is asking for; when does john live at home

OpenStudy (perl):

well im not so concerned with the exclusive or, since we said earlier that x can't be both -2 and -3 at the same time anyway

OpenStudy (perl):

' mr. smith likes apple pie or mr. smith likes spinach' . we dont know which one

OpenStudy (perl):

all we can say is that mr. smith likes at least one of {apple pie, spinach }

OpenStudy (perl):

possibly both

OpenStudy (amistre64):

if (x=a or x=b), then (f(x)=0) if (f(x) !=0 ) then (x!=a and x!=b) has the same truth value since its a contrapositive

OpenStudy (perl):

the way you are using it, still has a problem.

OpenStudy (amistre64):

if (x=a or x=b), then (f(x)=0) has the same truth value as if (f(x)=0), then (x=a or x=b) which means that we can construct an iff statment, biconditional

OpenStudy (perl):

' mr. smith likes apple pie or mr. smith likes spinach' . we dont know which one

OpenStudy (amistre64):

your misusing the phraseology again :)

OpenStudy (perl):

' x = -3 or x = -2' , we dont know which one

OpenStudy (amistre64):

if (smith eats apples or smith eats spinach); THEN __________ if (smith eats apples or smith eats spinach); then (smith is eating supper)

OpenStudy (amistre64):

propositions are the ors ands right?

OpenStudy (perl):

we want to deduce that x = { -3, -2} , we cannot deduce that from x = -2 or x = -3

jimthompson5910 (jim_thompson5910):

Another way to look at it "{m,n} is the solution set to x^2 + 5x + 6 = 0" ---> True (all quadratics have 2 solutions) "{m,n} is the solution set to (x+3)(x+2) = 0" ---> True ((x+3)(x+2) is equivalent to x^2 + 5x + 6) "{m,n} is the solution set to x+3=0 or x+2 = 0" ---> True (zero product property) "{m,n} is the solution set to x = -3 or x = -2" ---> True (Additive property of equality) If you work backwards (then forward again), you can say "{-3, -2} is the solution set to x^2 + 5x + 6 = 0" ---> True (all quadratics have 2 solutions) "{-3, -2} is the solution set to (x+3)(x+2) = 0" ---> True ((x+3)(x+2) is equivalent to x^2 + 5x + 6) "{-3, -2} is the solution set to x+3=0 or x+2 = 0" ---> True (zero product property) "{-3, -2} is the solution set to x = -3 or x = -2" ---> True (Additive property of equality) "{-3, -2} is the solution set to -3 = -3 or -2 = -2" ---> True (substution and reflexive property)